OpenStudy (anonymous):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation.

OpenStudy (anonymous):

WTF?

OpenStudy (anonymous):

What???

OpenStudy (anonymous):

Just bored, sorry i don't know the answer i'll go find my friend to help you out he's kinda smarty.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

@undeadknight26 can you help me?

undeadknight26 (undeadknight26):

Sorry i know this but its not coming to me right now...

OpenStudy (anonymous):

Ok that is ok, do you know anyone that can help me?

undeadknight26 (undeadknight26):

@SolomonZelman @hartnn @amistre64

OpenStudy (anonymous):

Thanks :)

OpenStudy (solomonzelman):

2x + 7y = 1 ---> times 3 ----> 6x +21y = 3 -3x – 4y = 5 ---> times 2 ----> -6x -8y = 10 ----------------- 0x +13y = 13 y=1 can you solve for x?

OpenStudy (anonymous):

well if y = 1 then shouldn't x = 0

OpenStudy (anonymous):

But all I needed to change was the first equation not both of them

OpenStudy (solomonzelman):

why would it be 0, no it's not zero, solve for x by plugging in 1 for y into one of the equations.

OpenStudy (anonymous):

Oh sorry I was looking at the very last equation on the bottom right. If y = 1 then x = -3

OpenStudy (solomonzelman):

See how I did it? now do 2x + 7(1) = 1 -3x – 4(1) = 5 2x+7=1 -3x-4=5 2x=-6 x=-3 -3x=9 x=-3 So y=1 and x=-3 is right.

OpenStudy (solomonzelman):

CLEAR?

OpenStudy (anonymous):

Kindof but i only need to change the first equation not both equations

OpenStudy (solomonzelman):

You can change both equations if you want to, in your case it's better to change both if you want to do elimination.

OpenStudy (anonymous):

But in the next part of the question it tells me to solve it using the new first equation and the origional second equation

OpenStudy (solomonzelman):

What?! What a poor school, I think they are referring to the "substitution" like this. 2x + 7y = 1 ---> x= (-7/2)y+(1/2) substitute (-7/2)x+(1/2) for y. into -3x – 4y = 5 -3x – 4[(-7/2)x+(1/2)] = 5 and solve for x. But why would they tell you to substitute it in this type of question?

OpenStudy (anonymous):

I am so sorry @SolomonZelman I had to go help wash my dogs, that is why I left. But now I am totally confused

OpenStudy (anonymous):

I haven't learned to do all that